[Brahms-l] Shift report 20050213 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 13 2005 - 23:58:21 EST
1. Started without beam. MCR had some problem in the beginning with the beam, however once solved the beam was good
2. Runs:
runno.  -  tr2  -  tr3
FS@4A2442, MRS@40A2255
r13171   -   8k  -   4k
r13172   - 198k  -  95k 
r13173   - 344k  - 153k
r13174   - 192k  - 161k

FS@4A3450, MRS@40A2255
r13175   - 151k  - 114k

non-TPC pedestal run: r13169, r13177 (and r13176, which however is incomplete)
TPC pedestal run: r13170, r1378

3. No detector problems to speak of.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 13 23:58:46 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 13 2005 - 23:59:00 EST